# Frenet Immersed Finite Element Spaces on Triangular Meshes

### v1

@@ 6.2. Numerical Experiments on $L^{2}$ Projection
In this experiment, we examine the approximation properties of the GC-IFE spaces through the convergence of the $L^{2}$ projection $\mathcal{P}_{h}u$. All three types of IFE constructions are tested. For polynomial degrees $m\leq 5$, the convergence results obtained with the GC-IFE basis by Reconstruction Approach 2 are presented in Figures 4, for both the $L^{2}$ norm and $H^{1}$ semi-norm. In this range of polynomial degrees, the general construction and Reconstruction Approach 1 perform comparably; therefore, their results are omitted for brevity. The observed convergence rates indicate that
$\|u-\mathcal{P}_{h}u\|_{L^{2}(\Omega)}\approx\mathcal{O}(h^{m+1}),\qquad|u-\mathcal{P}_{h}u|_{H^{1}(\Omega)}\approx\mathcal{O}(h^{m}),$
which are optional regarding the underlying polynomial space $\mathbb{P}_{m}$.
As reported in Section 6.1, for GC-IFE basis constructed with higher degree polynomials (e.g. $m=7$), the conditioning of the mass matrices differs significantly among the three bases. As shown in Figure 5, the accuracy of the approximation with the GC-IFE basis by the general construction deteriorates dramatically from the expected optimal order, whereas both reconstructed bases continue to exhibit good convergence. In particular, the GC-IFE basis by Reconstruction Approach 2 is more robust and maintains the optimal convergence rates in both the $L^{2}$ norm and $H^{1}$ semi-norm.
Figure 4. Convergence of $L^{2}$ projection error in $L^{2}$ and $H^{1}$ norms with the GC-IFE basis by Reconstruction Approach 2 for degree $1\leq m\leq 5$.
@@ 6.3. IFE-DG Solution for Elliptic Interface Problems
Next, we apply our GC-IFE spaces in the standard [B.Riviere_bk2008] symmetric interior penalty discontinuous Galerkin (SIPDG) method for solving the elliptic interface problems (1.1)-(1.5). Specifically, this method is to find $u_{h}\in{\mathcal{V}}^{m}_{\beta}({\mathcal{T}_{h}})$ such that
$\begin{split}\sum_{K\in\mathcal{T}_{h}}\int_{K}\beta\nabla u_{h}\cdot\nabla v_{h}d\mathbf{x}-\sum_{e\in\mathcal{E}^{i}_{h}}\int_{e}\{\beta\nabla u_{h}\cdot\mathbf{n}\}\left\llbracket v_{h}\right\rrbracket ds&-\sum_{e\in\mathcal{E}^{i}_{h}}\int_{e}\{\beta\nabla v_{h}\cdot\mathbf{n}\}\left\llbracket u_{h}\right\rrbracket ds\\ +\sum_{e\in\mathcal{E}^{i}_{h}}\frac{\sigma_{e}}{|e|}\int_{e}\left\llbracket u_{h}\right\rrbracket\,\left\llbracket v_{h}\right\rrbracket ds&=\int_{\Omega}fv_{h}d\mathbf{x},\qquad\forall v_{h}\in{\mathcal{V}}^{m}_{\beta}({\mathcal{T}_{h}}).\end{split}$ (6.52)
@@
$\sigma_{e}=\sigma_{0}\frac{\max(\beta^{+},\beta^{-})^{2}}{\min(\beta^{+},\beta^{-})},\quad\sigma_{0}=m^{2}.$
Figure 6 demonstrates the convergence of the DG-IFE solution $u_{h}$ to the interface problem by GC-IFE basis via the general construction. The performance of reconstructed basis $\{\lambda_{j}^{(1)}\}$ and $\{\lambda_{j}^{(2)}\}$ is very similar and is therefore omitted. These numerical results demonstrate that the DG IFE solutions produced with all three bases achieve the expected convergence orders in both the $L^{2}$ norm and $H^{1}$ norm for low polynomial degrees ($m\leq 4$).
For higher-degree cases, e.g., $p\geq 5$, the convergence behavior of the DG-IFE solution on a finer mesh, i.e., $h_{max}=1/60$, will deteriorate. Since the magnitude of the error in numerical solutions is around $10^{-12}$ or below, we think this behavior is attributed to finite-precision arithmetic and round-off errors. Hence, we report the convergence of the DG solution obtained with GC-IFE bases on coarser meshes in Figures 6 and 7. Overall, our numerical experiments indicate that the reconstructed bases exhibit more robust performance compared with the general construction.
Figure 6. Convergence of DG-IFE errors in $L^{2}$ and $H^{1}$ norm with GC-IFE basis by general construction for degree $1\leq m\leq 4$.
Figure 7. Convergence of DG-IFE errors in $L^{2}$ and $H^{1}$ norm with GC-IFE basis by Reconstruction Approach 2 for degrees 55 and $6$.
@@ 7. Conclusion
We have developed geometry-conforming immersed finite element (GC-IFE) spaces on triangular meshes for elliptic interface problems using a Frenet–Serret transformation that straightens curved interfaces and enforces jump conditions exactly. Building on prior work on rectangular meshes [2024AdjeridLinMeghaichi, 2025AdjeridLinMeghaichi2], we presented three construction procedures for higher-degree Frenet-IFE spaces, including reconstructed approaches that substantially improve the conditioning of the mass matrix.
Numerical experiments confirm the optimal approximation properties and robustness of the proposed GC-IFE spaces on triangular meshes. When employed in the standard interior penalty discontinuous Galerkin framework, the resulting scheme can achieve optimal convergence in both the $H^{1}$ and $L^{2}$ norms. The methodology developed here provides a foundation for future research on solving interface problems with higher-degree polynomials on general domains using interface-independent meshes.

### v2

@@ 6.2. Numerical Experiments on $L^{2}$ Projection
In this experiment, we examine the approximation properties of the GC-IFE spaces through the convergence of the $L^{2}$ projection $\mathcal{P}_{h}u$. All three types of IFE constructions are tested. For polynomial degrees $m\leq 3$, the convergence results obtained using Reconstruction Approach 2 are presented in Figure 4, in both the $L^{2}$ and semi-$H^{1}$-norms. For higher polynomial degrees $4\leq m\leq 7$, we use coarser meshes to minimize the impact from round-off errors. The convergence results using Reconstruction Approach 2 are presented in Figure 5. In both cases, the behaviors of the general construction and Reconstruction Approach 1 are very similar, and therefore their results are omitted for brevity. The observed convergence rates indicate that
$\|u-\mathcal{P}_{h}u\|_{L^{2}(\Omega)}\approx\mathcal{O}(h^{m+1}),\qquad|u-\mathcal{P}_{h}u|_{H^{1}(\Omega)}\approx\mathcal{O}(h^{m}).$
As discussed in Section 6.1, for higher-order approximations (e.g. $m=7$), the conditioning of the mass matrices differs significantly among the three bases. As shown in Figure 6, the accuracy of the approximation based on the general construction deteriorates significantly, whereas both reconstructed approaches continue to exhibit convergence.
Figure 4. Convergence of $L^{2}$ projection in $L^{2}$ and $H^{1}$ norms using Reconstruction Approach 2 with degree $1\leq m\leq 3$.
Figure 5. Convergence of $L^{2}$ projection in $L^{2}$ and $H^{1}$ norms using Reconstruction Approach 2 with degree $4\leq m\leq 7$.
@@ 6.3. IFE-DG Solution for Elliptic Interface Problems
Next, we apply our GC-IFE spaces in the symmetric interior penalty discontinuous Galerkin (SIPDG) method for solving the elliptic interface problems (1.1)-(1.5). Find $u_{h}\in{\mathcal{V}}^{m}_{\beta}({\mathcal{T}_{h}})$ such that
$\begin{split}\sum_{K\in\mathcal{T}_{h}}\int_{K}\beta\nabla u_{h}\cdot\nabla v_{h}d\mathbf{x}-\sum_{e\in\mathcal{E}^{i}_{h}}\int_{e}\{\beta\nabla u_{h}\cdot\mathbf{n}\}\left\llbracket v_{h}\right\rrbracket ds&\\ -\sum_{e\in\mathcal{E}^{i}_{h}}\int_{e}\{\beta\nabla v_{h}\cdot\mathbf{n}\}\left\llbracket u_{h}\right\rrbracket ds+\sum_{e\in\mathcal{E}^{i}_{h}}\frac{\sigma_{e}}{|e|}\int_{e}\left\llbracket u_{h}\right\rrbracket\,\left\llbracket v_{h}\right\rrbracket ds&=\int_{\Omega}fv_{h}d\mathbf{x},~\forall v_{h}\in{\mathcal{V}}^{m}_{\beta}({\mathcal{T}_{h}}).\end{split}$
@@
$\sigma_{e}=\sigma_{0}\frac{\beta_{\max}^{2}}{\beta_{\min}},\quad\sigma_{0}={\color[rgb]{0,0,0}m^{2}},$
where $\beta_{\max}=\max\{\beta^{+},\beta^{-}\}$ and $\beta_{\min}=\min\{\beta^{+},\beta^{-}\}$. In Figures 7 and 8, we present the convergence of SIPDG solutions using IFE bases from Reconstruction Approach 2. The performance of IFE basis from Reconstruction Approach 1 is very similar and is therefore omitted. These numerical results demonstrate that the DG solutions constructed with all three bases achieve the expected convergence orders in both the $L^{2}$ and $H^{1}$ norms.
Figure 7. Convergence of SIPDG solution in $L^{2}$ and $H^{1}$ norms using Reconstruction Approach 2 with degrees up to 3.
Figure 8. Convergence of SIPDG solution in $L^{2}$ and $H^{1}$ norms using Reconstruction Approach 2 with degrees $4\leq m\leq 7$.
The computational results using general construction are also similar for low-order polynomials ($m\leq 6$). For higher-degree approximation (e.g. $m=7$), we report the convergence of the DG solutions generated from all three bases in Figure 9. Again, the reconstructed bases exhibit more robust performance compared with the general construction.
@@
6.4. Non-rectangular Domain
In this example, we consider an interface problem defined on a non-rectangular domain. Since rectangular meshes are not applicable in this setting, we use triangular meshes to partition the domain. The computational domain is a regular hexagon with vertices
$A_{k}=\left(\cos\frac{(2k+1)\pi}{6},\ \sin\frac{(2k+1)\pi}{6}\right),\qquad 0\leq k\leq 5.$
The interface $\Gamma$ is taken to be the same circle as in previous examples, and the exact solution is defined as in (6.1). The triangular meshes are generated with MATLAB PDE Toolbox independent of the interface, as illustrated in the right plot of Figure 3.
The convergence of the $L^{2}$ projection using Reconstruction Approach 2 is reported in Figure 10. Optimal convergence rates are observed in both the $L^{2}$ norm and the semi-$H^{1}$ norm. The convergence of the SIPDG solution error is reported in Figure 11, where optimal decay rates are again achieved in both norms. The convergence behavior for higher-contrast coefficients is qualitatively similar and is therefore omitted. Likewise, for higher-degree approximations ($m\geq 5$), the results are analogous to those in Figure 5 and Figure 9 and are omitted for brevity.
Figure 10. $L^{2}$ Projection errors in $L^{2}$ and $H^{1}$ norms on hexagonal domain using Reconstruction Approach 2 with degrees up to $4$.
Figure 11. SIPDG errors in $L^{2}$ and $H^{1}$ norms on hexagonal domain using Reconstruction Approach 2 with degrees up to $4$.
@@ 7. Conclusion
We have developed geometry-conforming immersed finite element spaces on triangular meshes for elliptic interface problems using a Frenet–Serret mapping that straightens curved interfaces and enforces the jump conditions exactly. Extending earlier work on rectangular meshes, we introduced three high-order Frenet-IFE constructions, including reconstruction approaches that significantly improve mass matrix conditioning. Numerical experiments confirm the optimal approximation properties and robustness of the proposed spaces. When combined with interior penalty discontinuous Galerkin methods, the resulting schemes achieve optimal convergence rates in both the $H^{1}$- and $L^{2}$- norms. These results provide a foundation for future work on extending the Frenet mapping to handle boundary value problems with curved boundaries and high-order fitted-mesh methods for interface problems.
Acknowledgments